11 November 2008 Lab Meeting Agenda
† Hussam's research report today.
- Your personal priorities have been emailed to you by me. Please take note.
† Flu R01
- Courier service! Sample transport set up.
- Plaque assay status -- plaques and everything!Yeah!
- Mark - status of fixture?
† SEPSIS Project - CIMIT Bird Flu
- Jeff Braman will visit in early Dec.
- Do cDNA prep. Spec, gels, send to Jeff B. Shipping?
- Mark- Control circuit?
- Bacteria Experiment - Changing the teflon filter.
- Pluronic results? - straight channels. Treated/non, pluronic better. NEED UPDATE FROM JD
- Substrates - JZ tried with more gold and do all SEM/sample testing.
† Valve Array/Valve building (FJ)
- Status from Mark? - valves in transit somewhere.
- Frank should talk with Alex, Mark, and Jaephil and make sure we are not duplicating efforts/making something that is too big for portability.
- Frank met with Mark on Friday, has concerns along the lines of the above comment (duplicating efforts etc.) that will hopefully be addressed during/after lab meeting.
† Fraunhofer: LOAC
- Primer/dimer issue. Talked with QQ. We do not appear to have this problem with our chips.
† Biointerfaces group
- Mask made. Chrome mask.
- Preliminary E. coli experiment is ongoing.
† CIMIT- Colson Grant
- Preliminary plan for IRB submission at Harvard in place. Target date 1/1/09
- STAR-CD Renewal - January ($2K)
- Cathie look up her flu assay specs.
- Improve the gel concentration 1.2% agarose. See if you can separate out primer dimers.
- Look into Melting temp analysis on the 7300 - can only do with Sonali's assay.
- Only 1/5 chips are good at the hot embossing step. The vaccum chuck will be checked and o-rings changed - helped a bit.
* Fixuture for PCR chip. The heater and the thermocouples would be on the fixture.
- MinCheol's designs. Done. Data looks the same as old design. Extension time not long enough.
- Will do extension design with lambda phage assay.
- Mincheol will work on DNA migration simulation after IMECE 2008.
- Cathie submitted One pager Whitepaper. Waiting for response.
† Senior project
- Megan has mask designs, JD will check, send out this week.
- Depth of channels, check with JD and Sonali, because Sonali will be able to tell you how much assay volume.
- Mark will help/learn SU8 step.
- Teddy doing off chip controls work. Will consult with Megan and JD to get interface design going. Software control, too..LabVIEW.
- Valves in the mail --- find them! (MARK)
† F31: Cochlea
- Made changes and gsve to Cathie.
† Silica Optimization (Lambda):
- De-inhibit Methyl Methacrylate (100ml), put in fridge, make new columns.
- Ongoing. More concentrations.
- Pico Green assay figured out.
- Ethanol precipitation protocol and preparation of sodium acetate. ASAP.